NOTICE of CMS GENERAL MEETING

The General Meeting of the Canadian Mathematical Society will take place at 5:30 pm on Sunday, December 13, 1998 in the Sir John A. MacDonald Room of the Holiday Inn Kingston Waterfront, One Princess Street, Kingston, Ontario.

The proposed agenda is as follows:

1. Adoption of the agenda.

2. Minutes of the previous Annual General Meeting (June 14, 1998) *

3. Matters Arising

4. President’s Report (a) Strategic Planning

(b) World Math Year 2000

5. Executive Director and Secretary’s Report

6. Treasurer’s Report • Operating Fund Budget 1999

7. Reports from Committees (a) Women in Mathematics

(b) Research

(c) Publications

(d) Nominating

(e) Mathematical Olympiads

(f) International Affairs

(g) Human Rights

(h) Government Policy

(i) Fund Raising

(j) Finance

(k) Electronic Services

(l) Education

8. Other Business

9. Adjournment

(*) As printed in the November 1998 issue of the CMS Notes, Volume 30, No.7, pp. 19–22. Additional Information will be available on site.

AVIS de L’ASSEMBLÉE GÉNÉRALE

L’assemblée générale annuelle de la Société mathématique du Canada aura lieu à 17h30 le dimanche 13 décembre dans la salle Sir John A. MacDonald du Holiday Inn Kingston Waterloo, 1 rue Princess, Kingston, Ontario.

Voici le projet d’ordre du jour:

1. Adoption de l’ordre du jour

2. Procès verbal de la dernière assemblée générale annuelle (14 juin 1998) *

3. Affaires découlant du procès verbal

4. Rapport du président (a) Plan stratǵique

(b) l’An 2000 - année mondiale des mathématiques

5. Rapport du directeur administratif et secrétaire

6. Rapport du trésorier • Budget 1999 du fonds d’opération

7. Rapports des comités (a) Femmes en mathématiques

(b) Recherche

(c) Publications

(d) Mises en candidature

(e) Olympiades mathématiques

(f) Affaires internationales

(g) Droits de la personne

(h) Politiques gouvernementales

(i) Collecte des fonds

(j) Finance

(k) Services électronique

(l) Éducation

8. Affaires nouvelles

9. Ajournement

(*) S.V.P. voir les Notes de la SMC, Volume 30, no. 7. pp. 17–22. De plus amples renseignements seront fournis sur place.